### v1

# NoiseBench: Benchmarking the Impact of Real Label Noise on Named Entity Recognition

Abstract
Available training data for named entity recognition (NER) often contains a significant percentage of incorrect labels for entity types and entity boundaries. Such label noise poses challenges for supervised learning and may significantly deteriorate model quality. To address this, prior work proposed various noise-robust learning approaches capable of learning from data with partially incorrect labels. These approaches are typically evaluated using simulated noise where the labels in a clean dataset are automatically corrupted. However, as we show in this paper, this leads to unrealistic noise that is far easier to handle than real noise caused by human error or semi-automatic annotation. To enable the study of the impact of various types of real noise, we introduce NoiseBench, an NER benchmark consisting of clean training data corrupted with 6 types of real noise, including expert errors, crowdsourcing errors, automatic annotation errors and LLM errors. We present an analysis that shows that real noise is significantly more challenging than simulated noise, and show that current state-of-the-art models for noise-robust learning fall far short of their theoretically achievable upper bound. We release NoiseBench to the research community [^1] https://github.com/elenamer/NoiseBench.
NoiseBench: Benchmarking the Impact of Real Label Noise
on Named Entity Recognition
Elena Merdjanovska1,2, Ansar Aynetdinov1, Alan Akbik1,2 1Humboldt University of Berlin 2Science of Intelligence {elena.merdjanovska, aynetdia, alan.akbik}@hu-berlin.de
1 Introduction
Named entity recognition (NER) is the task of detecting and classifying named entities in text, such as the names of organizations or locations. All current state-of-the-art approaches for NER require supervision in the form of labeled training data, i.e. sentences in which named entities are marked and assigned their correct type. However, prior work found that available datasets for NER and other supervised tasks are affected by label noise, meaning that a certain percentage of entity labels are incorrect. For instance, the common NER dataset CoNLL-03 Tjong Kim Sang and De Meulder (2003) was estimated in various studies to have noise shares of between 5 and 7% Wang et al. (2019); Reiss et al. (2020); Rücker and Akbik (2023). Other NER datasets have also been found to contain a share of incorrect labels, with Ontonotes4 estimated around 8% and WNUT-17 around 18% Wang et al. (2019); Huang et al. (2021).
(a) Example of a partial match error induced by noise. Real noise makes a plausible mistake by labeling "UN" as ORG (organization), whereas simulated noise implausibly caused "Council" to be labeled.
(b) Examples of type and non-entity errors induced by noise. Real noise makes a plausible mistake by labeling "Swiss" as a LOC (location), whereas simulated noise implausibly labels "Grand Prix" as LOC. Real noise makes a plausible non-entity mistake by labeling "Sunday", whereas simulated noise labels "cycling".
(c) Example of a missing mention induced by noise. Real noise causes a plausible omission ("Olympic"), whereas simulated noise omits a trivial entity annotation ("Zurich")
Figure 1: Examples of text snippets with correct labels (top row) and two types of noise: Real noise from crowdsourcing (middle row) and simulated class-dependent noise (bottom row). This introduces different types of errors: (a) partial matches of correct entity mentions, (b) a wrong type and a non-entity mention and (c) a missing entity. We qualitatively find real noise to be more plausible than simulated noise.
Label noise introduces inconsistencies during training, which may significantly deteriorate model quality Zhang et al. (2021a). To address this issue, prior work proposed approaches for noise-robust learning aimed at mitigating the negative effects of the noisy training signal Song et al. (2022). However, the evaluation of these approaches has two main limitations.
Limitation 1: Simulated noise is too easy. Most current research in noise-robust learning relies on experiments with simulated label noise Tänzer et al. (2022); Klie et al. (2023). While this allows for evaluation in a controlled setting, it has been shown that simulated noise, even though it can model noise well to some extent, is much easier for deep learning models to disregard than real label noise Jiang et al. (2020); Zhu et al. (2022).
Refer to Figure 1 for an illustrative comparison between real and simulated noise for three example sentences, including different types of errors that occur in NER datasets. These examples demonstrate that simulated noise can introduce similar errors as real noise, however the choice of spans to mislabel is random and as a result often less plausible. This means that an approach shown to be robust to simulated noise may not in fact be robust to real noise in practice.
Limitation 2: Distinct types of real noise. Additionally, there exist many possible sources of "real" noise. For instance, expert labelers may make different mistakes than crowd workers Frenay and Verleysen (2014). Next to human labeling, there are widely-used automatic approaches to create NER-labeled datasets such as distant supervision from a knowledge base Mintz et al. (2009); Hedderich et al. (2021) and weak supervision using rules Zhang et al. (2021b). Lastly, current research investigates the use of LLMs to label datasets Golde et al. (2023); Wang et al. (2023). ](option1) choose one image of Figure 2, put the others in appendix
We postulate that these types of real noise differ in their characteristics, meaning that a noise-robust learning approach shown to perform well on one type of noise may not perform well on another. For this reason, we argue there is a need for evaluating noise-robustness across multiple label noise types.
Contributions. With this paper, we present NoiseBench, a new benchmark for measuring the impact of label noise in the training data on the prediction quality of trained NER models. In more detail, our contributions are:
  • We construct a noisy training dataset in 7 different variants, where each noisy variant contains the same sentences and is affected by one class of real errors, spanning errors made by experts, crowd workers, distant supervision, weak supervision and teacher LLMs.
  • We present a set of experiments that empirically show that real noise from NoiseBench is significantly more difficult for current approaches. We further find that during training, real noise is memorized immediately, whereas memorization of simulated noise is delayed.
  • We comparatively evaluate current state-of-the-art approaches for noise-robust learning on NoiseBench, and experimentally establish theoretical upper bounds.
Our analysis finds that no single current approach works best for all types of real noise, and that all current approaches fall far short of their theoretical upper bound. To enable the research community to leverage our benchmark in their evaluations, we publicly release all data and implementation.
inline]strengthen contributions. also make it clear that simulating noise is what everyone else is doing and creating an ideal NER noise model is also challenging.
2 NoiseBench
Our benchmark is derived from a subset of the classic CoNLL-03 dataset for NER in English, annotated with entities belonging to four classes. We chose this dataset since it has been extensively studied in the field, allowing us to integrate various prior works.
NoiseBench consists of the following parts: (1) A noise-free test split to evaluate trained models. (2) Seven variants of the training split, where six are annotated with different types of noise and one is without noise. The quality of the six noisy variants, in reference to the noise-free dataset, is presented in Table 2.
The training split contains 5,885 sentences from 400 documents, covering 9,685 entity mentions. The test split contains 3,427 sentences from 231 documents, covering 5,725 entity mentions.
#Entities %Errors
Noisy train split %Noise F1token F1entity Total Correct Missing (FN) Non-entity (FP) Type Partial
Expert 5.5 99.0 94.5 9,644 9,129 10.0 2.8 74.0 13.3
Crowd++ 15.3 96.7 84.7 8,607 7,751 59.6 8.7 17.0 14.7
Crowd 36.6 92.3 63.4 7,188 5,352 61.9 10.2 16.0 11.9
Distant 31.3 92.9 68.7 7,329 5,846 65.4 10.5 12.9 11.1
Weak 40.4 91.9 59.6 10,640 6,058 17.4 34.6 36.3 11.8
LLM 45.6 87.4 54.4 11,349 5,726 22.5 45.4 28.3 3.7
Table 1: Overview of the noisy training splits in NoiseBench. The table shows the noise level, the micro-averaged token-level F1 score (F1 token ), micro-averaged entity-level F1 (F1 entity ), the number of entities (Total), number of correct entities (Correct) and share of each error type: missing mentions (Missing (FN)), labelled non-entity mentions (Non-entity (FN)) and wrong type (Type), as well as the number of partial matches (Partial). All metrics are in comparison to the Clean split.
2.1 Types of Noise
In the following, we discuss each training split and the type of noise it models.
2.1.1 Noise-Free Data
Our benchmark requires two splits without any label noise: A clean test split to evaluate models trained on noisy training data, and a Clean training split to measure the upper bound performance.
Since the original annotations of CoNLL-03 have been shown to be noisy Wang et al. (2019); Reiss et al. (2020), we use the labels of CleanCoNLL (Rücker and Akbik, 2023), a recently released resource in which 7% of all original annotations were semi-automatically relabeled. In their evaluation, Rücker and Akbik (2023) find their resulting dataset to be of very high quality and largely improved consistency. The Clean Test split in our benchmark is the standard CoNLL-03 test split, with the CleanCoNLL labels.
2.1.2 Expert Errors
The largest portion of machine learning datasets relies on manual annotation by domain experts to provide high-quality labels. However, errors have been found to occur even in expert annotation, affecting even well-known benchmarks, though usually with relatively low noise shares of under 10% Northcutt et al. (2021b); Song et al. (2022). To represent such noise, our benchmark includes a variant of the train split called Expert, which contains the original CoNLL-03 annotations. As Table 1 shows, this split has a noise share of 5.5% and is thus the split with lowest noise.
2.1.3 Crowdsourcing Errors
In addition to expert annotations, labeled datasets can be obtained via crowdsourcing. These annotations are heavily prone to human annotation errors, which can happen for different reasons (Frenay and Verleysen, 2014). In order to create noisy variants of the train set in our benchmark representing real-world human errors, we utilize the crowdsourced labels by Rodrigues et al. (2014). This study involves 47 crowd workers labelling a subset of the CoNLL-03 dataset, of around 400 news articles. They released their dataset and all annotations produced by each crowd worker. We selected only the sentences where the tokenization matched the Clean variant, resulting in 5,885 sentences.
We include two noisy training splits based on crowd annotations into our benchmark: (1) In the first, Crowd, we do a simple majority vote over all annotations provided for each token, i.e. the baseline method for aggregating crowdsourced annotations. (2) In the second, Crowd++, we use an oracle version of the majority vote, selected by either taking the correct label if it is provided by any of the annotators or, in the absence of a correct label, by choosing the label with the majority of votes. This version represents the upper bound of crowdsourced labels given a perfect label aggregation method. As Table 1 shows, the noise share of Crowd (36.6%) is considerably higher than Crowd++ (15.3%).
2.1.4 Distant Supervision
One approach for labeling data without human participation is distant supervision Mintz et al. (2009), where entity mentions in target datasets are matched to entity types in knowledge bases (KBs). The matching can be achieved by simple string matching, use of regular expressions or heuristics.
We include a Distant noisy training variant in our benchmark, adapted from the annotations by Liang et al. (2020) [^2] Available under Apache 2.0 license that use the Wikidata corpus and gazetteers collected from multiple online sources as external knowledge bases. After initial POS tagging, the unlabeled sentences were matched with the knowledge bases.
This process results in incomplete annotations due to limited coverage over entity types of KBs. This explains the rather high number of missing entities and the overall noise level (31.3%) of the Distant training variant, as shown in Table 1.
2.1.5 Weak Supervision
Another approach aimed at reducing manual annotation efforts is weak supervision. Here, labels are obtained using a number of ’weak’ supervision sources, such as heuristics or expression-based rules. Each weak source is typically specialised to detect only a subset of the correct labels.
We use the labels from the approach by Lison et al. (2020)2 to create our Weak label set. This covers 16 weak labeling sources, those used in Zhang et al. (2021b), including heuristics, gazetteers and predictions of NER models trained on other corpora. An example heuristic is detecting PER (person) entities using a pre-defined list of first names.
We aggregate the weak label sets with simple majority voting. We apply majority vote on every token with at least one entity label assigned to it, following Zhang et al. (2021b). Due to the large number of labelling sources, majority voting yields a large number of entities, as shown in Table 1, including many false positives. As a result, the Weak label set has a high noise share of 40.4%.
2.1.6 LLM Teacher Models
Our benchmark includes a noisy variant of the train split annotated by an LLM. This follows recent efforts that use LLMs for dataset generation Wang et al. (2023). Here, the main idea is to pass a natural language description of the annotation task and target classes to an LLM, and provide sentences that the LLM should label. Current works find that LLMs are able to generate high quality labels for some tasks (e.g. sentiment classification) while for others (e.g. NER and question type categorization) the resulting labels are very noisy.
We created the LLM variant using the Fabricator toolkit Golde et al. (2023) by prompting GPT3.5 for named entities in our training dataset. To use LLM outputs for annotation of NER datasets, a certain output format is required. To achieve this, we provide one example with the correct output format in each prompt. This example is the same for each sentence we wish to annotate, which we refer to as a static one-shot setting. The example sentence was selected from the remainder of the CoNLL-03 training split, which consists of all sentences not included in our benchmark.
As Table 1 shows, the LLM label set results in the highest noise share of 45.6%. This is mainly due to the large number of nouns incorrectly identified as entity mentions, which also makes this the label set with the largest number of entity annotations out of the variants in NoiseBench.
2.2 Statistics
An overview of NoiseBench is given in Table 1. The table shows token-level F1 score and entity-level F1 score expressed as percentages. We define the noise level (%Noise) in terms of the entity-level F1 score, as $100-\%F1$. The noise levels of the noisy splits range from 5.5 to 45.6 percent.
Furthermore, the table shows the total number of entities, the number of correct entities, as well as the share of different error types. The errors are categorized into 4 main categories: missing mentions, non-entity mentions (false positives), incorrect entity type (where the boundary is correct, but type incorrect) and partial matches. Partial matches are special cases where the type is correct, however the mention boundary is only partially correct. Refer to Figure 1 for examples.
We observe that the Crowd++, Crowd and Distant label sets have a lower total number of entity annotations than the Clean dataset, and the largest portion of errors are missing mentions. Conversely, the Weak and LLM label sets have more annotations than the Clean dataset, and most of the errors are either an incorrect mention or incorrect type. Most of the errors in the Expert label set are due to incorrect type. Regarding the number of partial matches, for almost all noise types, they make up between 10% and 15% of all errors.
3 Comparing Real and Simulated Noise
We first use NoiseBench to investigate how real label noise affects NER model performance in comparison to simulated noise. For this, we conduct two experiments: the first one addresses the impact of each type of training noise on the clean test set performance, and the second one compares training dynamics under real and simulated label noise to highlight the differences in noise memorization.
Real noise Simulated noise $\Delta$
%Noise F1 %Noise F1 F1
Clean 0 94.0 ±0.0 - - -
Expert 5.5 89.8±0.2 5.9 93.7±0.2 3.9
Crowd++ 15.3 86.7±0.3 17.9 88.9±0.4 2.2
Crowd 36.6 70.5±0.6 41.3 72.4±1.0 1.8
Distant 31.3 70.8±0.1 39.2 74.5±0.4 3.7
Weak 40.4 65.9±0.4 41.2 63.1±0.8 -2.8
LLM 45.6 62.6±0.4 47.2 68.6±1.3 6.0
Average 74.4±0.3 76.9±0.7 2.5
Table 2: F1 scores on the Clean Test split of the baseline FLERT approach, fine-tuned on different noisy variants of the training set. The scores are averages of 3 runs. The column $\Delta$ (difference) refers to the difference in F1 score on the test split when training on a dataset with real noise compared to simulated class-dependent noise.
3.1 Noise Simulation Methods
We consider two noise simulation methods, namely the simple uniform noise used in most prior work and a more involved oracle class-dependent noise method that we design to mirror each noisy variant in NoiseBench.
Uniform noise. Uniform noise corrupts samples into any other label with a uniform probability distribution, given a target noise share. Studies investigating simulated noise in the NER task commonly rely on variants of this method Mayhew et al. (2019); Tänzer et al. (2022).
Oracle class-dependent noise. Class-dependent noise is based on the knowledge that some pairs of classes are more likely to be mislabeled than others. It is defined by a noise transition matrix, which contains the mislabeling probabilities between all pairs of classes Hedderich et al. (2021). We design an oracle version of class-dependent noise, where the per-class mislabeling probabilities of real noise are known. This allows us to investigate class-dependent noise in an ideal case, where it is able to mirror real noise closely, even though this is not possible in practice. This method mirrors real noise by utilizing the token-level mislabeling frequencies as probabilities to form a noise transition matrix.
Using each noise simulation method, we created 6 label sets, corresponding to each noise level in NoiseBench. It should be noted that the simulated labels replicate the token-level F1 scores of the real noisy labels, however the entity-level F1 and sentence-level accuracy can deviate.
3.2 Experimental Setup
In both experiments, we train a baseline approach for NER on each noisy variant of the training split, as well as on the additional simulated noise.
Validation splits. We evaluate the setting in which all available data to train a model is noisy, including the validation set.
To obtain noisy validation sets for each of our 7 dataset variants, we split the noisy datasets into training and validation sets. All sentences from 66 news documents from 1996-08-24 comprise the validation set, which is around 17% of all sentences, and are left out from model training and used for hyperparameter tuning.
Baseline. For NER, as a baseline approach, we fine-tune an xlm-roberta-large transformer using the FLERT approach Schweter and Akbik (2021). It improves upon the regular fine-tuning setup by considering document-level features of a sentence to be tagged. We use a learning rate of 5e-6 and a batch size of 32, for a fixed number of 10 epochs. These parameters were obtained according to the performance on a noisy validation set, keeping in mind that larger batch sizes are more robust to noise Rolnick et al. (2017). As an evaluation metric we use entity-level micro F1-score.
Expert Crowd++ Distant
Real
(a) Real Expert  - 5.5% noise
(b) Real Crowd++  - 18% noise
(c) Real Distant  - 31.3% noise
Simulated
(d) Simul. Expert  - 6% noise
(e) Simul. Crowd++  - 15% noise
(f) Simul. Distant  - 39% noise
Figure 2: Comparison of model performance during extended training. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated (class-dependent) noisy label set. The graphs show the F1 score on the noisy training labels and on the clean training labels, for 3 noise types. The plots are averages of 3 runs. The remaining 3 noise types are shown in Appendix C.
3.3 Experiment 1: Impact of Label Noise on Test Performance
In the first experiment, we compare how the clean test set performance is impacted by the 6 types of real label noise when present in the training set. In addition, we provide the same comparison for corresponding simulated noisy label sets.
3.3.1 Results and Discussion
The results for uniform noise are shown in Appendix A. We initially established that uniform noise is significantly less challenging for the model, so in the results from Experiments 1 and 2 we chose to focus our analysis solely on oracle class-dependent noise.
The main results from Experiment 1 for oracle class-dependent noise are shown in Table 2. Additional analysis of the results can be found in Appendix B. Following are our main observations.
Label noise degrades performance. When we compare the test F1 scores of the real noisy variants with the average score of 93.99 achieved when training on the Clean variant, we can see that model performance is affected by each noise type. In all cases, the noisy training variants yield lower test scores, even when training with the low-noise Expert variant. As the noise levels increase, the impact on model performance becomes more pronounced. This shows that the baseline model lacks robustness to any of the real noise types. When we compare the test F1 scores of the simulated noisy variants, we can see that noise of 5.9 in the training set does not hurt model performance and results in a score comparable to training on the Clean variant. However, as simulated noise levels increase, they do degrade the prediction scores on the test set.
Real noise is more difficult. Furthermore, when we compare the real noisy label sets with their equivalent simulated noisy variants, we can observe that the simulated training variants show a score of around 2.5 percentage points higher on average than the real label sets. This shows that for predictive NER models, real noise is more difficult to overcome than simulated noise. In other words, models are more likely to overfit to real noisy labels, rather than simulated ones.
3.4 Experiment 2: Memorization of Noise
Prior analysis has found that there are distinct phases of learning when training a model on data with label noise Arpit et al. (2017). This has been referred to as a generalization phase, where models learn patterns that generalize well to clean data, followed by a memorization phase, where models overfit to the label noise and deteriorate in prediction quality Tänzer et al. (2022).
To investigate this phenomenon for real and simulated noise, we extend the training stage to 100 epochs. At the end of each epoch, we measure the F1 score of the model on both the noisy training split it is being trained on, and separately on the clean training split. The difference between these two scores allows us to measure memorization.
3.4.1 Results and Discussion
In Table 2 we show training curves from training with real and simulated variants of NoiseBench  for 3 noise types: Expert, Crowd++ and Distant. We plot two scores: the F1-score on the respective noisy variant of the training set, and the F1 score on the Clean variant of the training set. In all training curves, we can observe the memorization effect, with each model perfectly fitting the noisy data by the end of training and reaching an F1 score close to 1.
Delayed memorization of simulated noise. However, we note that with simulated noise (see Table 4, 4, 4) this happens much later in the training process than with real noise. In addition, the training curves of simulated noise show a stage during the early epochs where the score on the clean labels is consistently higher than the score on the noisy labels. This confirms previous findings that the model is able to learn general patterns first, before starting to memorize the noise.
Immediate memorization of real noise. With real noise this does not happen and the model starts fitting the noisy labels from the beginning of training (see Table 4, 4, 4). As a result, the score on the clean labels is consistently lower than the score on the noisy labels, during the entire training run.
Our experiments find that real noise does not display distinct generalization/memorization phases during training, and rather immediately begins with memorization. This makes intuitive sense, as real noise has underlying patterns that may be extracted during learning. This lends further evidence to the increased challenges and the need to evaluate noise-robust learning with real noise.
4 Evaluating Noise-Robust Learning
Having established the difficulty of real noise, we now use NoiseBench to perform a comparative evaluation of widely-used noise-robust learning approaches. Our goal is to determine their effectiveness in the presence of real label noise, and to establish upper bounds of what noise-robust learning could ideally achieve.
4.1 Compared Approaches
We surveyed current state-of-the-art methods for noise-robust NER and found that many approaches rely on the same underlying ideas for handling label noise. In the following, we group approaches by the underlying idea, select a state-of-the-art representative for each group and, if possible, derive an upper bound method for each noise-robust method. For more details about the implementation of compared approaches refer to Appendix D.
4.1.1 Learning from a Clean Subset
The first family of approaches relies on utilizing the subset of each noisy dataset in which all labels are correct. They either filter out all likely incorrect annotations and learn only from a clean subset Northcutt et al. (2021a). Or they derive confidence weights for each annotation so that annotations judged to be of higher quality feature more during training, as in the CrossWeigh and COSINE approaches Wang et al. (2019); Yu et al. (2021). As representative of this class of approaches, we chose confident learning Northcutt et al. (2021a).
Upper bound: Oracle subset. To obtain an upper bound for this family of approaches, we use an oracle to select the subset of clean sentences from each of the noisy training splits in NoiseBench. We then use a standard fine-tuning approach only on this subset. This setting illustrates the best-case scenario for training on a clean subset only.
4.1.2 Delaying Memorization
Another family of noise-robust learning approaches seeks to leverage the two phases of learning (generalization and memorization) we discussed in Section 3.4. They seek to either draw our the generalization phase or cease training before memorization begins. While our experiments indicate that these two phases do not exist for real noise, we nevertheless include this family of approaches in our evaluation since they are widely used and to gain additional insights into their behavior on real noise. As representative of this class of approaches, we chose co-regularization (Zhou and Chen, 2021).
Upper bound: Oracle stopping. To obtain an upper bound for this family of approaches, we use a simple stopping criterion based on the score on the clean test set at the end of each epoch. We use the epoch of best generalization to report the final score. This simulates an ideal stopping (albeit at the granularity of full epochs).
Clean Expert Crowd++ Crowd Distant Weak LLM Avg.
Baseline 93.99±0.04 89.84±0.19 86.71±0.29 70.52±0.62 70.75±0.13 65.87±0.36 62.60±0.39 77.18
Upper bounds
Oracle subset - 90.31±0.28 91.83±0.33 85.95±0.59 83.07±0.59 81.13±1.25 75.70±1.10 85.99
Oracle stopping 94.06±0.07 89.88±0.19 87.23±0.24 71.04±0.88 71.84±0.61 66.98±0.22 63.64±0.44 77.81
Noise-robust learning
Co-regularization 93.65±0.11 89.57±0.28 87.11±0.32 72.67±0.87 72.89±0.28 66.82±0.15 62.23±0.76 77.85
Confident learning 93.71±0.31 90.01±0.15 86.53±0.23 69.99±0.97 71.41±0.34 65.81±0.46 61.75±0.56 77.03
BOND 92.46±0.48 88.97±0.16 86.13±0.81 73.33±1.00 73.67±1.15 65.72±0.74 61.46±0.65 77.39
Table 3: Performance of noise-robust approaches on the Clean Test set, when training on NoiseBench training split variants. Results are expressed in terms of F1 score. Each score is averaged over 3 runs.
4.1.3 Combined Approaches
While the approach discussed so far each build on the individual ideas of identifying a clean subset or delaying memorization, many current approaches in fact combine multiple of such ideas in multi-stage pipelines Liang et al. (2020); Yu et al. (2021); Wang et al. (2022). As representative of such approaches, we evaluate BOND Liang et al. (2020), which combines pseudo-labeling in a student-teacher setup and sample selection.
No upper bound for pseudo-labeling. We cannot derive a separate upper bound for pseudo-labeling, as the best case scenario here would mean that all noisy labels are replaced by correct labels. This means that the upper bound for pseudo-labeling is the same as training on fully clean data.
4.2 Results
Table 3 summarizes the evaluation results. We make the following observations:
Clean subset has high potential. The upper bound of training only on clean examples is the best achievable result for all types of noise. This was to be expected, since in this scenario the labels are not contaminated by any noise. This indicates that noise-robust approaches focusing on identifying the clean subset have high potential. Oracle stopping, on the other hand, does not achieve the same level of performance as the oracle subset, while only slightly outperforming the FLERT baseline. This is in line with our findings in Experiment 2 that the early-learning generalization phase is skipped when training with real noise, indicating that early stopping approaches have less potential.
Small benefit of noise-robust approaches. Evidently, there is no single best approach for all noise types. While for all noise types some of the noise-robust approaches are able to outperform the baseline, they do so only slightly, with the exception of BOND, on the Distant and Crowd variants. Still, their performance is far below the upper bound. This raises the issue of trade-offs of existing noise-robust learning approaches, since they often require additional hyperparameter tuning or incur computational costs, but only lead to slight improvements over the baseline in the presence of real noise.
5 Related work
There are a few benchmarks for learning with label noise and related topics. The WRENCH benchmark Zhang et al. (2021b) focuses only on weak supervision labels for multiple tasks, with the emphasis on combining multiple weak labelling sources. Klie et al. (2023) compare a large number of methods for the detection of annotation errors. Multiple tasks are included, including NER on CoNLL-03, where they evaluate the detection of expert errors, concluding that most approaches are not successful at this. Similarly, Chong et al. (2022) evaluate annotation error detection on datasets with noise only from crowdsourced labels, for part-of-speech tagging and natural language inference tasks. Liu et al. (2022) propose a benchmark for text classification under label noise, where they re-annotate an existing sentiment classification dataset and construct noisy label sets according to annotator disagreements; however, they do not publish these label sets. NoisyWikiHow, a benchmark for intention identification has also been presented Wu et al. (2023), where the authors propose a method to simulate realistic noise that closely imitates human errors by producing heterogeneous and instance-dependent errors. For NER in Estonian, Hedderich et al. (2021) introduce the NoisyNER, which includes multiple noise levels obtained from distant supervision approaches with varying quality.
6 Conclusion
In this paper, we address the issue of label noise in the NER task. We introduce a new benchmark, based on the commonly used NER dataset CoNLL-03 in English, for evaluating the impact of 6 distinct types of real label noise on the same set of sentences, with varying degrees of difficulty.
We demonstrate that real noise causes transformer-based language models to immediately memorize the noise pattern, making real label noise a more challenging problem than simulated label noise, even in the case of oracle class-dependent noise informed by the characteristics of real noise.
We further investigate this by evaluating popular approaches aimed at mitigating this issue and showcasing their limitations. Our experiments indicate that current methods fall far short of what can be potentially achieved on almost all of the noisy datasets in NoiseBench, leaving room for improvement. We hope that NoiseBench aids other researchers in the further development of more effective noise-robust approaches and in characterizing the differences between types of real noise.
Limitations
We identify two limitations in our study.
First, it focuses on the scenario when the entire available dataset could be noisy and we do not have access to a small, high-quality labelled, data subset. While this is a certainly scenario which reflects a large number of real-world cases, it could be argued that in some situations it is realistic to have the resources to ensure a subset of the data is clean, with high-quality annotations. However, when this is the case, Zhu et al. (2023) showcased that this clean data would be better utilized by directly fine-tuning the models on it, instead of using it for validation. Therefore, we argue that this alternative setup is not particularly useful for the evaluation of label-noise-robust approaches.
Second, this benchmark is limited to the English language, because it is dependent on the availability of various labelling sources for the widely used CoNLL dataset. Learning under label noise is a greater challenge in domains where data is limited, which is why investigating this problem for low-resource languages would be beneficial for the research community. Nevertheless, we believe the findings regarding noise robustness obtained with our benchmark would be transferable to data in other languages.
Acknowledgements
Elena Merdjanovska and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. Ansar Aynetdinov and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Emmy Noether grant “Eidetic Representations of Natural Language” (project number 448414230).
Appendix A Baseline for uniform noise
In Table 4 the results from Experiment 1 for uniform noise are shown. We can see the the model is quite robust to uniform noise and that it results in higher test performance, when compared to real noise of the same level. The average difference in F1 scores in 17 percentage points, which is why we focus on the more realistic oracle class-dependent noise simulation method in the remainder of the results of Experiments 1 and 2.
Real noise Uniform noise $\Delta$
%Noise F1 %Noise F1 F1
Clean 0 94.0 ±0.0 - - -
Expert 5.5 89.8±0.2 5.4 93.8±0.3 4
Crowd++ 15.3 86.7±0.3 16.1 92.8 ±0.5 6.1
Crowd 36.6 70.5±0.6 36.7 88.4 ±0.3 17.9
Distant 31.3 70.8±0.1 31.7 90.0 ±0.5 19.3
Weak 40.4 65.9±0.4 42 91.7 ±0.2 25.8
LLM 45.6 62.6±0.4 47.3 89.7 ±0.6 27.1
Average 74.4±0.3 91.40.4 17
Table 4: F1 scores on the Clean Test split of the baseline FLERT approach, fine-tuned on different noisy variants of the training set. The scores are averages of 3 runs. The column $\Delta$ (difference) refers to the difference in F1 score on the test split when training on a dataset with real noise compared to uniform noise.
Appendix B Generalization to unseen entities
(a) Real noise
(b) Simulated class-dependent noise
Figure 3: Comparison of F1 scores on different subsets of entities in the test set: all, seen (clean), seen (noisy) and unseen.
We evaluate the generalization of the baseline model from Experiment 1 (Section 3.3) by calculating F1 scores for seen and unseen entities separately, and further distinguishing seen entities according to whether their label in the training set was clean or noise-corrupted. On Figure 3(a) we observe that for the Expert and Crowd++ noise types, the performance for the seen (clean) entities and the unseen entities is comparable, which indicates the model has the ability to generalize to unseen entities well. The overall drop in performance when training with these noisy training splits is largely due to the low performance on the seen (noisy) entities. However, this detrimental effect of memorization is limited to the seen errors and does not affect the performance on unseen entities much, for noise levels up to 15%. As for the remaining training splits with noise levels of 30% and above, we can see that noise affects the performance on unseen entities as well. Furthermore, on Figure 3(b), we see that for simulated noise, for Expert and Crowd++ , the performance on seen noisy entities is close to the performance on seen clean entities. This means that at low noise levels, the models are able to disregard the noisy patterns and predict the same entities correctly when they appear in the test set, instead of memorizing the noisy label.
Appendix C Memorization of Crowd, Weak and LLM training variants
The memorization plots from Experiment 2, as an extension of Section 3.4 for the Crowd, Weak and LLM dataset variants are shown in Table 4
Crowd Weak LLM
Real
(a) Real Crowd  - 36.6% noise
(b) Real Weak  - 40.4% noise
(c) Real LLM  - 44.6% noise
Simulated
(d) Simul. Crowd  - 41.3% noise
(e) Simul. Weak  - 41.2% noise
(f) Simul. LLM  - 47.2% noise
Figure 4: Comparison of model performance during extended training. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated noisy label set (class-dependent noise). The graphs show both the F1 score on the noisy training labels and on the clean training labels, for 3 different noise types. The plots are averages of 3 runs.
. We again observe immediate memorization of real noise and delayed memorization of simulated noise.
Appendix D Implementation details
D.1 Confident learning
We use regular transformer fine-tuning and obtain predicted probabilities for each sample in the training dataset using cross-validation. The number of folds is the only parameter in this approach and we performed a small search before choosing 3 folds. We use the implementation by Klie et al. (2023) to adapt this approach for NER by aggregating token-level predictions. We perform the final sample selection on the sentence level, training the model only using sentences that do not contain entities flagged as errors or have missing entities.
D.2 Co-regularization
We use two classifiers and perform a hyperparameter sweep as suggested by the authors.
D.3 BOND
Since we cannot perform early stopping on a clean validation set, which is suggested for both stages by the authors, we rely on the findings reported by Tänzer et al. (2022) and stop the first stage after the first 4 epochs. The second stage is limited to 6 epochs in order to reproduce the same training length as in the FLERT baseline. We adopt the hyperparameters reported as best by the authors for the CoNLL-03 dataset, and use hard pseudo-labels in the second stage.