### v2

# NoiseBench: Benchmarking the Impact of Real Label Noise on Named Entity Recognition

Abstract
Available training data for named entity recognition (NER) often contains a significant percentage of incorrect labels for entity types and entity boundaries. Such label noise poses challenges for supervised learning and may significantly deteriorate model quality. To address this, prior work proposed various noise-robust learning approaches capable of learning from data with partially incorrect labels. These approaches are typically evaluated using simulated noise where the labels in a clean dataset are automatically corrupted. However, as we show in this paper, this leads to unrealistic noise that is far easier to handle than real noise caused by human error or semi-automatic annotation. To enable the study of the impact of various types of real noise, we introduce NoiseBench, an NER benchmark consisting of clean training data corrupted with 6 types of real noise, including expert errors, crowdsourcing errors, automatic annotation errors and LLM errors. We present an analysis that shows that real noise is significantly more challenging than simulated noise, and show that current state-of-the-art models for noise-robust learning fall far short of their achievable upper bound. We release NoiseBench for both English and German to the research community [^1] https://github.com/elenamer/NoiseBench.
NoiseBench: Benchmarking the Impact of Real Label Noise
on Named Entity Recognition
Elena Merdjanovska1,2, Ansar Aynetdinov1, Alan Akbik1,2 1Humboldt-Universität zu Berlin 2Science of Intelligence {elena.merdjanovska, aynetdia, alan.akbik}@hu-berlin.de
1 Introduction
Named entity recognition (NER) is the task of detecting and classifying named entities in text, such as the names of organizations or locations. Current state-of-the-art approaches for NER still require supervision in the form of labeled training data Zaratiana et al. (2023), i.e. sentences in which named entities are marked and assigned their correct type. However, prior work found that available datasets for NER and other supervised tasks are affected by label noise, meaning that a certain percentage of entity labels are incorrect. For instance, the common NER dataset CoNLL-03 Tjong Kim Sang and De Meulder (2003) was estimated in various studies to have noise shares of between 5 and 7% Wang et al. (2019); Reiss et al. (2020); Rücker and Akbik (2023). Other NER datasets have also been found to contain a share of incorrect labels, with OntoNotes4 estimated around 8% and WNUT-17 around 18% Wang et al. (2019); Huang et al. (2021).
(a) Example of a partial match error induced by noise. Real noise makes a plausible mistake by labeling "UN" as ORG (organization), whereas simulated noise implausibly caused "Council" to be labeled.
(b) Examples of type and non-entity errors induced by noise. Real noise makes a plausible mistake by labeling "Swiss" as a LOC (location), whereas simulated noise implausibly labels "Grand Prix" as LOC. Real noise makes a plausible non-entity mistake by labeling "Sunday", whereas simulated noise labels "cycling".
(c) Example of a missing mention induced by noise. Real noise causes a plausible omission ("Olympic"), whereas simulated noise omits a trivial entity annotation ("Zurich")
Figure 1: Examples of text snippets with correct labels (top row) and two types of noise: Real noise from crowdsourcing (middle row) and simulated class-dependent noise (bottom row). This introduces different types of errors: (a) partial matches of correct entity mentions, (b) a wrong type and a non-entity mention and (c) a missing entity. We qualitatively find real noise to be more plausible than simulated noise.
Label noise introduces inconsistencies during training, which may significantly deteriorate model quality Zhang et al. (2021a). To address this issue, prior work proposed approaches for noise-robust learning aimed at mitigating the negative effects of the noisy training signal Song et al. (2022). However, the evaluation of these approaches has two main limitations.
]introduction seems a bit long?
Limitation 1: Simulated noise is too easy. Most current research in noise-robust learning relies on experiments with simulated label noise Tänzer et al. (2022); Klie et al. (2023). While this allows for evaluation in a controlled setting, it has been shown that simulated noise, even though it can model noise well to some extent, is much easier for deep learning models to disregard than real label noise Jiang et al. (2020); Zhu et al. (2022).
Refer to Figure 1 for an illustrative comparison between real and simulated noise for three example sentences, including different types of errors that occur in NER datasets. These examples demonstrate that simulated noise can introduce similar errors as real noise, however the choice of spans to mislabel is random and as a result often less plausible. This means that an approach shown to be robust to simulated noise may not in fact be robust to real noise in practice.
Limitation 2: Distinct types of real noise. Additionally, there exist many possible sources of "real" noise. For instance, expert labelers may make different mistakes than crowd workers Frenay and Verleysen (2014). Next to human labeling, there are widely-used automatic approaches to create NER-labeled datasets such as distant supervision from a knowledge base Mintz et al. (2009); Hedderich et al. (2021) and weak supervision using rules Zhang et al. (2021b). Lastly, current research investigates the use of LLMs to label datasets Golde et al. (2023); Wang et al. (2023). ](option1) choose one image of Figure 2, put the others in appendix
We postulate that these types of real noise differ in their characteristics, meaning that a noise-robust learning approach shown to perform well on one type of noise may not perform well on another. For this reason, we argue there is a need for evaluating noise-robustness across multiple label noise types.
Contributions. With this paper, we present NoiseBench, a new benchmark for measuring the impact of label noise in the training data on the prediction quality of trained NER models. In more detail, our contributions are:
  • We construct a noisy training dataset in 7 different variants, where each noisy variant contains the same sentences and is affected by one class of real errors, spanning errors made by experts, crowd workers, distant supervision, weak supervision and teacher LLMs.
  • We present a set of experiments that empirically show that real noise from NoiseBench is significantly more difficult for current approaches. We further find that during training, real noise is memorized immediately, whereas memorization of simulated noise is delayed.
  • We comparatively evaluate current state-of-the-art approaches for noise-robust learning on NoiseBench, and experimentally establish upper bounds.
Our analysis finds that no single current approach works best for all types of real noise, and that all current approaches fall far short of their theoretical upper bound. To enable the research community to leverage our benchmark in their evaluations, we publicly release all data and implementation.
inline]strengthen contributions. also make it clear that simulating noise is what everyone else is doing and creating an ideal NER noise model is also challenging.
2 NoiseBench
Our benchmark is derived from a subset of the classic CoNLL-03 dataset for NER in English, annotated with entities belonging to four classes. We chose this dataset since it has been extensively studied in the field, allowing us to integrate various prior works. We derive a similar benchmark for NER in German, in Section 5 and Appendix A.
NoiseBench consists of the following parts: (1) A noise-free test split to evaluate trained models. (2) Seven variants of the training split, where six are annotated with different types of noise and one is without noise. Table 1 presents the quality of the six noisy variants w.r.t. the noise-free dataset.
The training split contains 5,885 sentences from 400 documents, covering 9,685 entity mentions. The test split contains 3,427 sentences from 231 documents, covering 5,725 entity mentions.
#Entities %Errors
Noisy train split %Noise F1token F1entity Total Correct Missing (FN) Non-entity (FP) Type Partial
Expert 5.5 99.0 94.5 9,644 9,129 10.0 2.8 74.0 13.3
Crowd++ 15.3 96.7 84.7 8,607 7,751 59.6 8.7 17.0 14.7
Crowd 36.6 92.3 63.4 7,188 5,352 61.9 10.2 16.0 11.9
Distant 31.3 92.9 68.7 7,329 5,846 65.4 10.5 12.9 11.1
Weak 40.4 91.9 59.6 10,640 6,058 17.4 34.6 36.3 11.8
LLM 45.6 87.4 54.4 11,349 5,726 22.5 45.4 28.3 3.7
Table 1: Overview of the noisy training splits in NoiseBench. The table shows the noise level, the micro-averaged token-level F1 score (F1token), micro-averaged entity-level F1 (F1entity), the number of entities (Total), number of correct entities (Correct) and share of each error type: missing mentions (Missing (FN)), non-entity mentions (Non-entity (FP)), wrong type (Type) and partial matches (Partial). All metrics are in reference to the Clean split.
2.1 Types of Noise
In the following, we discuss each training split and the type of noise it models.
2.1.1 Noise-Free Data
Our benchmark requires two splits without any label noise: A clean test split to evaluate models trained on noisy training data, and a Clean training split to measure the upper bound performance.
Since the original annotations of CoNLL-03 have been shown to be noisy Wang et al. (2019); Reiss et al. (2020), we use the labels of CleanCoNLL (Rücker and Akbik, 2023), a recently released resource in which 7% of all original annotations were semi-automatically relabeled. In their evaluation, Rücker and Akbik (2023) find their resulting dataset to be of very high quality and largely improved consistency. The Clean Test split in our benchmark is the standard CoNLL-03 test split, with the CleanCoNLL labels.
2.1.2 Expert Errors
Most machine learning datasets are created using manual annotation by domain experts that provide high-quality labels. However, errors have been found to occur even in expert annotation, affecting even well-known benchmarks, though usually with relatively low noise shares of under 10% Northcutt et al. (2021b); Song et al. (2022). To represent such noise, our benchmark includes a variant of the train split called Expert, which contains the original CoNLL-03 annotations. As Table 1 shows, this split has a noise share of 5.5% and is thus the split with lowest noise.
2.1.3 Crowdsourcing Errors
Crowdsourcing is a less costly alternative to expert annotation, but also more prone to annotation errors (Frenay and Verleysen, 2014). In order to create noisy variants of the train set representing real-world human errors, we utilize the crowdsourced labels by Rodrigues et al. (2014). This study involves 47 crowd workers labelling a subset of the English CoNLL-03 dataset, of around 400 news articles. They released their dataset and all annotations produced by each crowd worker. We selected only the sentences where the tokenization matched the Clean variant, resulting in 5,885 sentences.
We include two noisy training splits based on crowd annotations into our benchmark: (1) In the first, Crowd, we do a simple majority vote over all annotations provided for each token, i.e. the baseline method for aggregating crowdsourced annotations. (2) In the second, Crowd++, we use an oracle version of the majority vote, selected by either taking the correct label if it is provided by any of the annotators or, in the absence of a correct label, by choosing the label with the majority of votes. This version represents the upper bound of crowdsourced labels given a perfect label aggregation method. As Table 1 shows, the noise share of Crowd (36.6%) is considerably higher than Crowd++ (15.3%).
2.1.4 Distant Supervision
One approach for labeling data without human participation is distant supervision Mintz et al. (2009), where entity mentions in target datasets are matched to entity types in knowledge bases (KBs).
We include a Distant noisy training variant in our benchmark, adapted from the annotations by Liang et al. (2020) [^2] Available under Apache 2.0 license that use the Wikidata corpus and gazetteers collected from multiple online sources as external knowledge bases. After initial POS tagging, the unlabeled sentences were matched with the knowledge bases. This process results in incomplete annotations due to limited coverage over entity types of KBs. This explains the rather high number of missing entities and the overall noise level (31.3%) of the Distant training variant, as shown in Table 1.
2.1.5 Weak Supervision
Another approach aimed at reducing manual annotation efforts is weak supervision. Here, labels are obtained using a number of “weak” supervision sources, such as heuristics or expression-based rules. Each weak source is typically specialized to detect only a subset of the correct labels.
We use the labels from the approach by Lison et al. (2020)2 to create our Weak label set. This covers 16 weak labeling sources Zhang et al. (2021b), including heuristics, gazetteers and predictions of NER models trained on other corpora. An example heuristic is detecting PER (person) entities using a pre-defined list of first names.
We aggregate the weak label sets with simple majority voting. We apply majority vote on every token with at least one entity label assigned to it, following Zhang et al. (2021b). Due to the large number of labelling sources, majority voting yields a large number of entities, as shown in Table 1, including many false positives. As a result, the Weak label set has a high noise share of 40.4%.
2.1.6 LLM Teacher Models
Our benchmark includes a noisy variant of the train split annotated by an LLM. This follows recent efforts that use LLMs for dataset generation Wang et al. (2023). Here, the main idea is to pass a description of the annotation task and target classes to an LLM, and provide sentences to label. LLMs are able to generate high quality labels for some tasks (e.g. sentiment classification) while for others (e.g. NER and question type categorization) the resulting labels are very noisy Golde et al. (2023).
We created the LLM variant using the Fabricator toolkit Golde et al. (2023) by prompting GPT3.5 for named entities in our training dataset. To use LLM outputs for annotation of NER datasets, a certain output format is required. To achieve this, we provide one example with the correct output format in each prompt. This example is the same for each sentence we wish to annotate, which we refer to as a static one-shot setting. The example sentence was selected from the remainder of the CoNLL-03 training split, which consists of all sentences not included in our benchmark.
As Table 1 shows, the LLM label set results in the highest noise share of 45.6%. This is mainly due to the large number of nouns incorrectly identified as entity mentions, which also makes this the label set with the largest number of entity annotations out of the variants in NoiseBench.
2.2 Statistics
An overview of NoiseBench is given in Table 1. The table shows token-level F1 score and entity-level F1 score expressed as percentages. We define the noise level (%Noise) in terms of the entity-level F1 score, as $100-\%F1$. The noise levels of the noisy splits range from 5.5 to 45.6 percent.
The table also shows the share of different error types. The errors are categorized into 4 main categories: missing mentions, non-entity mentions (false positives), incorrect entity type (where the boundary is correct, but type incorrect) and partial matches. Partial matches are special cases where the type is correct, but the mention boundary is only partially correct. Refer to Figure 1 for examples.
We observe that the Crowd++, Crowd and Distant label sets have a lower total number of entity annotations than the Clean dataset, and the largest portion of errors are missing mentions. Conversely, the Weak and LLM label sets have more annotations than the Clean dataset, and most of the errors are either an incorrect mention or incorrect type. Most of the errors in the Expert label set are due to incorrect type. Regarding the number of partial matches, for almost all noise types, they make up between 10% and 15% of all errors.
3 Comparing Real and Simulated Noise
We first use NoiseBench to investigate how real label noise affects NER model performance in comparison to simulated noise. For this, we conduct two experiments: the first one addresses the impact of each type of training noise on the clean test set performance, and the second one compares training dynamics under real and simulated label noise to highlight the differences in noise memorization.
3.1 Noise Simulation Methods
We consider two noise simulation methods, namely the simple uniform noise used in most prior work and a more involved oracle class-dependent noise method that we design to mirror each noisy variant in NoiseBench.
Uniform noise. Uniform noise corrupts samples into any other label with a uniform probability distribution, given a target noise share. Studies investigating simulated noise in the NER task commonly rely on variants of this method Mayhew et al. (2019); Tänzer et al. (2022).
Oracle class-dependent noise. Class-dependent noise is based on the knowledge that some pairs of classes are more likely to be mislabeled than others. It is defined by a noise transition matrix, which contains the mislabeling probabilities between all pairs of classes Hedderich et al. (2021). We design an oracle version of class-dependent noise, where the per-class mislabeling probabilities of real noise are known. This allows us to investigate class-dependent noise in an ideal case, where it is able to mirror real noise closely, even though this is not possible in practice. This method mirrors real noise by utilizing the token-level mislabeling frequencies as probabilities to form a noise transition matrix.
Using each noise simulation method, we created 6 label sets, corresponding to each noise level in NoiseBench. It should be noted that the simulated labels replicate the token-level F1 scores of the real noisy labels, however the entity-level F1 and sentence-level accuracy can deviate.
3.2 Experimental Setup
In both experiments, we train a baseline approach for NER on each noisy variant of the training split, as well as on the additional simulated noise.
Validation splits. We evaluate the setting in which all available data to train a model is noisy, including the validation set. To obtain noisy validation sets for each of our 7 dataset variants, we split the noisy datasets into training and validation sets. All sentences from 66 news documents from 1996-08-24 comprise the validation set, which is around 17% of all sentences, and are left out from model training and used for hyperparameter tuning.
Baseline. For NER, as a baseline approach, we fine-tune an xlm-roberta-large transformer using the FLERT approach Schweter and Akbik (2021). It improves upon the regular fine-tuning setup by considering document-level features of a sentence to be tagged. We use a learning rate of 5e-6 and a batch size of 32, for a fixed number of 10 epochs, without early stopping. These parameters were obtained according to the performance on a noisy validation set, keeping in mind that larger batch sizes are more robust to noise Rolnick et al. (2017). We use entity-level micro F1-score.
3.3 Experiment 1: Impact of Label Noise on Test Performance
In the first experiment, we compare how the clean test set performance is impacted by the 6 types of real label noise when present in the training set. In addition, we provide the same comparison for corresponding simulated noisy label sets.
3.3.1 Results and Discussion
The results for uniform noise are shown in Appendix D. We initially established that uniform noise is less challenging for the model, so in the results for Experiments 1 and 2 we chose to focus solely on oracle class-dependent noise.
The main results from Experiment 1 for oracle class-dependent noise are shown in Table 2. Following are our main observations.
Label noise degrades performance. When we compare the test F1 scores of the real noisy variants with the average score of 93.99 achieved when training on the Clean variant, we see that model performance is affected by each noise type. As the noise levels increase, this impact becomes more pronounced, showing that the baseline model lacks robustness to real noise. Comparing the test F1 scores of the simulated noisy variants, we can see that noise of 5.9% in the training set results in a score comparable to training on the Clean variant. However, as simulated noise levels increase, the noise does degrade test set scores.
Real noise Simulated noise $\Delta$
%Noise F1 %Noise F1 F1
Clean 0 94.0 ±0.0 - - -
Expert 5.5 89.8±0.2 5.9 93.7±0.2 3.9
Crowd++ 15.3 86.7±0.3 17.9 88.9±0.4 2.2
Crowd 36.6 70.5±0.6 41.3 72.4±1.0 1.8
Distant 31.3 70.8±0.1 39.2 74.5±0.4 3.7
Weak 40.4 65.9±0.4 41.2 63.1±0.8 -2.8
LLM 45.6 62.6±0.4 47.2 68.6±1.3 6.0
Average 74.4±0.3 76.9±0.7 2.5
Table 2: F1 scores on the Clean Test split of the baseline FLERT approach, fine-tuned on different noisy variants of the training set. The scores are averages of 3 runs. The column $\Delta$ (difference) refers to the difference in F1 score on the test split when training on a dataset with real noise compared to simulated class-dependent noise.
Real noise is more difficult. Furthermore, when we compare the real noisy label sets with their equivalent simulated noisy variants, we can observe that the simulated training variants show a score of around 2.5 percentage points higher on average than the real label sets. This shows that for predictive NER models, real noise is more difficult to overcome than simulated noise. In other words, models are more likely to overfit to real noisy labels, rather than simulated ones.
(a) Real noise
(b) Simulated class-dependent noise
Figure 2: F1 scores on different subsets of entities in the test set: all, seen (clean), seen (noisy) and unseen.
Models generalize to unseen entities well. Figure 2 shows F1 scores for seen and unseen entities separately, further distinguishing seen entities by whether their label in the training set was clean or noise-corrupted. In Figure 2(a) we see that for the Expert and Crowd++ noise types, the score on the seen (clean) and the unseen entities is comparable, which indicates the model has the ability to generalize to unseen entities well. As for the remaining training splits with noise levels of over 30%, noise also affects the performance on unseen entities.
Simulated noisy patterns are disregarded. For all real noise types, the score on the seen (noisy) entities is low. With simulated noise however, in Figure 2(b) we see that for Expert and Crowd++, the score on the seen-noisy entities and seen-clean entities is close. This means that at low noise levels, the models are able to disregard simulated noisy patterns and predict the same entities correctly when they appear in the test set.
Expert Crowd++ Distant
Real
(a) Real Expert  - 5.5% noise
(b) Real Crowd++  - 18% noise
(c) Real Distant  - 31.3% noise
Simulated
(d) Simul. Expert  - 6% noise
(e) Simul. Crowd++  - 15% noise
(f) Simul. Distant  - 39% noise
Figure 3: Comparison of model performance during extended training. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on corresponding simulated (class-dependent) noisy labels. The plots are averages of 3 runs. The graphs for Crowd, Weak and LLM are shown in Appendix E.
3.4 Experiment 2: Memorization of Noise
Prior analysis has found that there are distinct phases of learning when training a model on data with label noise Arpit et al. (2017). This has been referred to as a generalization phase, where models learn patterns that generalize well to clean data, followed by a memorization phase, where models overfit to the label noise and deteriorate in prediction quality Tänzer et al. (2022).
To investigate this phenomenon for real and simulated noise, we extend the training stage to 100 epochs. At the end of each epoch, we measure the F1 score of the model on both the noisy training split it is being trained on, and separately on the clean training split. The difference between these two scores allows us to measure memorization.
3.4.1 Results and Discussion
In Figure 3 we show training curves from training with real and simulated variants of NoiseBench  for 3 noise types: Expert, Crowd++ and Distant. We plot two scores: the F1-score on the respective noisy variant of the training set, and the F1 score on the Clean variant of the training set. In all training curves, we can observe the memorization effect, with each model perfectly fitting the noisy data by the end of training and reaching an F1 score close to 1.
Delayed memorization of simulated noise. However, we note that with simulated noise (see Figure 3, 3, 3) this happens much later in the training process than with real noise. In addition, the training curves of simulated noise show a stage during the early epochs where the score on the clean labels is consistently higher than the score on the noisy labels. This confirms previous findings that the model is able to learn general patterns first, before starting to memorize the noise.
Immediate memorization of real noise. With real noise, this does not happen and the model starts fitting the noisy labels from the beginning (see Figure 3, 3, 3). As a result, the score on the clean labels is consistently lower than the score on the noisy labels, throughout the training run [^3] We confirm this finding for German in Appendix A.2.2..
Our experiments find that real noise does not display distinct generalization/memorization phases during training, and rather immediately begins with memorization [^4] We confirm this finding for a smaller model, as well as a randomly initialized model in Appendix E.. This makes intuitive sense, as real noise has underlying patterns that may be extracted during learning. This lends further evidence to the increased challenges and the need to evaluate noise-robust learning with real noise.
4 Evaluating Noise-Robust Learning
Having established the difficulty of real noise, we now use NoiseBench to perform a comparative evaluation of widely-used noise-robust learning approaches. Our goal is to determine their effectiveness in the presence of real label noise, and to establish upper bounds of what noise-robust learning could ideally achieve.
4.1 Compared Approaches
We surveyed current state-of-the-art methods for noise-robust NER and found that many approaches rely on the same underlying ideas for handling label noise. In the following, we group approaches by the underlying idea, select a state-of-the-art representative for each group and, if possible, derive an upper bound method for each group. For more details about the implementation of compared approaches refer to Appendix B.
4.1.1 Learning from a Clean Subset
The first family of approaches relies on utilizing the subset of each noisy dataset in which all labels are correct. One type of these approaches filters out all likely incorrect annotations and learns only from a clean subset. Another type derives confidence weights for each sample so that annotations judged to be of higher quality feature more during training.
As a representative of the former type of approaches targeting clean subsets of noisy datasets, we chose Confident Learning Northcutt et al. (2021a), while the latter type is represented by CrossWeigh Wang et al. (2019) and Learn-To-Reweight (L2R) Ren et al. (2018).
Upper bound: Oracle subset. To obtain an upper bound for this family of approaches, we use an oracle to select the subset of clean sentences from each of the noisy training splits in NoiseBench. We then use the baseline fine-tuning approach only on this subset, illustrating a best-case scenario.
4.1.2 Delaying Memorization
Another family of noise-robust learning approaches seeks to leverage the two phases of learning (generalization and memorization) we discussed in Section 3.4. They seek to either draw our the generalization phase or cease training before memorization begins. While our experiments indicate that these two phases do not exist for real noise, we nevertheless include this family of approaches in our evaluation since they are widely used. As representative of this class of approaches, we chose co-regularization (Zhou and Chen, 2021).
Upper bound: Oracle stopping. To obtain an upper bound for this family of approaches, we use a simple stopping criterion based on the score on the clean test set at the end of each epoch. We use the epoch of best generalization to report the final score. This simulates an ideal stopping.
Clean Expert Crowd++ Crowd Distant Weak LLM Avg.
Baseline 93.99±0.04 89.84±0.19 86.71±0.29 70.52±0.62 70.75±0.13 65.87±0.36 62.60±0.39 77.18
Upper bounds
Oracle subset - 90.31±0.28 91.83±0.33 85.95±0.59 83.07±0.59 81.13±1.25 75.70±1.10 85.99
Oracle stopping 94.06±0.07 89.88±0.19 87.23±0.24 71.04±0.88 71.84±0.61 66.98±0.22 63.64±0.44 77.81
Additional clean data 94.14±0.18 90.04±0.30 89.14±0.67 81.70±0.95 80.19±0.73 71.66±1.66 72.06±1.32 82.70
Noise-robust learning
Confident learning 93.71±0.31 90.01±0.15 86.53±0.23 69.99±0.97 71.41±0.34 65.81±0.46 61.75±0.56 77.03
CrossWeigh 93.50±0.12 89.68±0.49 85.01±0.83 64.95±1.18 70.55±0.24 66.15±0.24 60.87±1.40 75.82
L2R 90.29±2.12 82.10±4.10 79.91±2.27 67.51±1.01 65.45±2.01 63.36±0.34 65.29±4.15 73.42
Co-regularization 93.65±0.11 89.55±0.22 86.91±0.31 72.22±0.73 70.45±0.22 65.52±0.62 62.23±0.76 77.22
BOND 89.92±0.71 86.78±0.35 86.13±0.81 74.12±0.49 73.62±0.70 66.60±0.36 60.99±0.77 76.88
MSR 92.83±0.16 89.53±0.48 88.45±1.08 68.44±3.79 75.80±1.41 69.48±0.32 64.57±1.22 78.44
Table 3: Performance of noise-robust approaches on the Clean test set, when training on NoiseBench training split variants. Results are expressed in terms of F1 score. Each score is averaged over 3 runs.
4.1.3 Combined Approaches
While the approaches discussed so far each build on the individual ideas of identifying a clean subset or delaying memorization, many current approaches in fact combine multiple of such ideas in multi-stage pipelines Liang et al. (2020); Yu et al. (2021); Wang et al. (2022). As representative of such approaches, we evaluate BOND Liang et al. (2020) and meta self-refinement (MSR) Zhu et al. (2023a), both of which combine pseudo-labeling in a student-teacher setup and confidence-based sample selection.
]moved some implementation details to appendix
No upper bound for pseudo-labeling. We cannot derive a separate upper bound for pseudo-labeling, as the best case scenario here would mean that all noisy labels are replaced by correct labels, which is the same as training on the Clean dataset.
4.1.4 Additional Clean Data
We include a further upper bound for the scenario in which a small amount of high quality noise-free data is available. This is inspired by the extensive analysis of the use of clean validation data in Zhu et al. (2023b). Here, after first training on the noisy training set, they use a small clean dataset to continue fine-tuning the model. We include this upper bound to measure the accuracy gains that may be achieved if one were to invest effort in manually annotating additional noise-free data.
4.2 Results
Table 3 summarizes the evaluation results. We make the following observations:
Identifying a clean subset has highest potential. The upper bound of training only the clean subset of each noisy split (see "Oracle subset" in Table 3) achieves the best scores of all upper bounds. This makes intuitive sense as training is performed only over fully clean sentences, albeit a smaller subset of the full training data as all noisy sentences are filtered out. Similarly, we find strong improvements for the "Additional clean data" upper bound. Oracle stopping, on the other hand, does not achieve the same level of performance as the oracle subset, only slightly outperforming the FLERT baseline. This is in line with our findings in Experiment 2 that the early-learning generalization phase is skipped when training with real noise. This indicates that noise-robust learning approaches that target early stopping have little potential.
Small benefit of noise-robust approaches. Evidently, there is no single best approach for all noise types. For each noise type, at least one noise-robust approach outperforms the baseline, however on average most of them are comparable to it. Only MSR outperforms the baseline averaged over all noise types, bringing improvements for Crowd++, Distant, Weak and LLM. Additionally, L2R works well for LLM noise and BOND for Crowd. Still, the performance is far below the upper bound. This raises the issue of trade-offs of existing noise-robust learning approaches, since they often require additional hyperparameter tuning or incur computational costs, but only lead to slight improvements in the presence of real noise.
5 Ablation: NoiseBench for German
Using the German sentences in CoNLL-03, we created a noisy label benchmark for German to confirm our findings for a different language. Following the English counterpart described in Section 2, it consists of (1) a noise-free test split to evaluate trained models and (2) three variants of the training split, where two are annotated with different types of noise and one is the ground truth. The two types of noise include Expert labels, with 16.2% noise and LLM labels, with 54% noise. More details can be found in Appendix A.1.
5.1 Experimental Results
Clean Expert LLM Avg.
Baseline 90.24±0.2 79.02±0.4 57.86±0.4 75.7
Upper bounds
Oracle subset - 83.11±0.6 82.72±0.6 82.9
Oracle stopping 90.50±0.2 79.48±0.3 61.81±0.9 77.3
Additional clean data 89.86±0.9 82.85±1.7 69.50±1.4 80.7
Noise-robust learning
Confident learning 90.00±0.3 79.57±0.3 58.03±0.2 75.9
CrossWeigh 90.11±0.3 78.32±0.3 57.50±0.8 75.3
L2R 81.45±3.1 74.14±0.8 53.07±1.6 69.6
Co-regularization 88.50±0.2 78.49±0.2 54.47±0.5 73.8
BOND 86.53±0.3 77.56±0.5 55.89±0.6 73.3
MSR 85.34±0.5 76.42±0.4 64.00±0.7 75.3
Table 4: German variant: Performance of noise-robust approaches on the Clean test set, when training on each training split variant. Results are expressed in terms of F1 score. Each score is averaged over 3 runs.
Table 4 shows the results of the noise-robust approaches and upper bounds when training on the German datasets. More experimental details and results can be found in Appendix A.2.
Oracle subset score reaches an upper limit. Regarding the upper bounds, we see that the performance of the oracle subset of Expert and LLM is close, meaning that the 4000555See Appendix C for the size of the oracle subset. clean sentences in the LLM subset are already enough to reach an F1 score over 82. Despite having more samples, the Expert subset does not result in a much higher score. This could signify that the remaining sentences, not included in the Expert subset, are difficult examples necessary to properly learn the task.
Poor performance of noise-robust approaches. Regarding the noise-robust approaches, only Confident Learning is able to match and slightly outperform the baseline. All other methods mainly perform poorly on the German dataset, even below the baseline, with the exception of the improvement brought by MSR on the LLM dataset.
6 Related Work
There are a few benchmarks for learning with label noise and related areas. The WRENCH benchmark Zhang et al. (2021b) focuses only on weak supervision labels for multiple tasks, with the emphasis on combining multiple weak labelling sources. Klie et al. (2023) compare a large number of methods for the detection of annotation errors. Multiple tasks are included, including NER on CoNLL-03, where they evaluate the detection of expert errors, concluding that most approaches are not successful at this. Similarly, Chong et al. (2022) evaluate annotation error detection on datasets with noise only from crowdsourced labels, for part-of-speech tagging and natural language inference tasks. Liu et al. (2022) propose a benchmark for text classification under label noise, where they re-annotate an existing sentiment classification dataset and construct noisy label sets according to annotator disagreements; however, they do not publish these label sets. NoisyWikiHow, a benchmark for intention identification has also been presented Wu et al. (2023), where the authors propose a method to simulate realistic noise that imitates human errors by producing heterogeneous and instance-dependent errors. For NER in Estonian, Hedderich et al. (2021) introduce the NoisyNER, which includes multiple noise levels obtained from distant supervision approaches with varying quality. MultiCoNERv2 Fetahu et al. (2023) addresses textual noise in the input data itself (e.g. typos), instead of label noise.
7 Conclusion
In this paper, we address the issue of label noise in the NER task. We introduce a new benchmark, based on the commonly used NER dataset CoNLL-03, for evaluating the impact of 6 distinct types of real label noise on the same set of sentences, with varying degrees of difficulty.
We demonstrated that real noise causes transformer-based language models to immediately memorize the noise pattern, making real label noise a more challenging problem than simulated label noise, even in the case of oracle class-dependent noise informed by the characteristics of real noise.
We further presented an evaluation of popular noise-robust learning approaches. Our experiments indicate that current methods fall far short of what can potentially be achieved on the noise types in NoiseBench and that approaches that focus on automatically identifying a clean subset of labels have the highest potential. We hope that NoiseBench aids other researchers in the further development of more effective noise-robust approaches.
Limitations
This paper focuses on the scenario when the entire available dataset could be noisy and we do not have access to a small, high-quality labelled, data subset. While this is a certainly scenario which reflects a large number of real-world cases, it could be argued that in some situations it is realistic to have the resources to ensure a subset of the data is clean, with high-quality annotations. However, when this is the case, Zhu et al. (2023b) showcased that this clean data would be better utilized by directly fine-tuning the models on it, instead of using it for validation. Therefore, we argue that this alternative setup is not particularly useful for the evaluation of label-noise-robust approaches.
Acknowledgements
We thank all reviewers for their valuable comments. Elena Merdjanovska and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. Ansar Aynetdinov and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Emmy Noether grant “Eidetic Representations of Natural Language” (project number 448414230).
Appendix A NoiseBench for German: Additional Details
Following are additional details about the German version of NoiseBench, as well as results from Experiment 2 for this dataset.
A.1 Overview
The training split contains 12,705 sentences from 553 documents, covering 10,008 entity mentions. The test split contains 3,160 sentences from 155 documents, covering 3,051 entity mentions.
A.1.1 Noise-Free Data
As noise-free labels for the German part of CoNLL-03, we take the updated annotations from 2006666More details about the revision of the labels can be found in the ner.tz file, downloaded from https://www.clips.uantwerpen.be/conll2003/ner/, more specifically in the /ner/etc.2006/revision.txt and ner/etc.2006/guide.pdf files.. This updated label set is considered ground-truth by the research community.
When compared to the original CoNLL-03 German labels, he most changes in this label set are for the MISC class, most notably the removal of adjectives derived from names. This can alternatively be considered an update in annotation guidelines. With this, it should be noted that for the German dataset we do not have access to labels with verified high quality, as we do for the English counterpart with the CleanCoNLL labels.
A.1.2 Expert Errors
Similar as for English, we take the original CoNLL-03 labels Tjong Kim Sang and De Meulder (2003) as labels with expert errors. This results in a noise share of 16.2%.
A.1.3 LLM Teacher Models
Similar as for English, we use GPT3.5 to create a noisy version of the training split annotated by an LLM. This results in a high noise share of 54%.
A.1.4 Statistics
When compared to the noise shares for Expert and LLM in NoiseBench in Table 1 (5.5% and 45.6%), the noise shares for German are higher (16.2% and 54% respectively). This is due to LLMs performing more poorly on languages other than English, as well as due to fewer research efforts focusing on re-annotating and cleaning the German part of CoNLL-03, resulting in less consistent labels.
We see that most of the errors in German are non-entity mentions. Overall, we also note that type errors and partial matches are much less prominent here (less than 10%), even though they formed a larger part of the errors in NoiseBench.
#Entities %Errors
Noisy train split %Noise F1token F1entity Total Correct Missing (FN) Non-entity (FP) Type Partial
Expert 16.2 98.2 83.8 11852 9156 14.5 73.0 6.1 6.4
LLM 54.0 92.1 46.0 16526 6102 19.8 69.9 3.2 7.0
Table 5: Overview of the noisy training splits in NoiseBench for German. The table shows the noise level, the micro-averaged token-level F1 score (F1token), micro-averaged entity-level F1 (F1entity), the number of entities (Total), number of correct entities (Correct) and share of each error type: missing mentions (Missing (FN)), non-entity mentions (Non-entity (FP)), wrong type (Type) and partial matches (Partial). All metrics are in comparison to the Clean split.
A.2 Experiments
A.2.1 Validation Split
We take the last 96 documents from the training split to serve as a validation set, corresponding to roughly 17% of all sentences.
A.2.2 Experiment 2
We performed Experiment 2 for the German version of NoiseBench, where the goal is to observe the memorization of label noise. The resulting graphs for both noise types are shown in Figure 4. We observe a similar behaviour as in the English part of NoiseBench, where the real noisy datasets are memorized immediately.
(a) Expert - 16.4% noise
(b) LLM - 54.7% noise
Figure 4: Comparison of model performance during extended training, for the German dataset.
Appendix B Implementation Details
In all our experiments with noise-robust methods we use the same xlm-roberta-large transformer as in the FLERT baseline with a batch size of 32, except for L2R, for which we used a batch size of 16 due to VRAM constraints.
B.1 Confident Learning
We use regular transformer fine-tuning and obtain predicted probabilities for each sample in the training dataset using cross-validation. The number of folds is the only parameter in this approach and we performed a small search before choosing 3 folds. We use the implementation by Klie et al. (2023) to adapt this approach for NER by aggregating token-level predictions. We perform the final sample selection on the sentence level, training the model only using sentences that do not contain entities flagged as errors or have missing entities.
B.2 Co-Regularization
We perform a hyperparameter sweep as suggested by the authors, and choose the best performing ones on the validation set of the respective noise type.
B.3 BOND
In our experiments we found that limiting BOND’s first stage of training to 1 epoch is not enough for optimal performance, hence why we rely on the findings reported by Tänzer et al. (2022) and stop the first stage after the first 3 epochs. The second stage is limited to 7 epochs in order to reproduce the same training length as in the FLERT baseline. We update the teacher model in the second stage of training every 2 epochs as suggested by the authors for the CoNLL-03 dataset, and use hard pseudo-labels in the second stage, which we found to outperform soft pseudo-labels in our experiments.
B.4 CrossWeigh
We ran the CrossWeigh framework with 5 folds and 3 iterations, because according to the ablation experiments ran by the authors, higher numbers did not bring significant performance improvements. For a fair comparison, we adjust the CrossWeigh framework to use transformer fine-tuning as a base model. For the final training run using the sample weights, we used the same FLERT approach as in the baseline.
B.5 L2R
We rely on the implementation provided by Zhu et al. (2023b), test out two meta-learning rates while keeping the model learning rate fixed at 5e-6, and perform the validation step every 0.1 epoch with a patience of 10 validation steps.
B.6 MSR
We used the implementation provided by the authors Zhu et al. (2023a) and the hyperparameters they selected for CoNLL, as stated in their paper. For the German dataset, we used xlm-roberta-base as a multilingual model.
B.7 Upper Bound: Additional Clean Data
This upper bound assumes an additional small dataset with high-quality labels is available. We fix this number to 100 sentences, which are randomly chosen from the validation split (otherwise not used to train the models). This training setting first fine-tunes the baseline model for 10 epochs, and then continues fine-tuning only on the small clean dataset for 5 more epochs.
B.8 GPT3.5
To obtain an LLM-annotated variant of the training splits, for NoiseBench we used gpt-3.5-turbo-0613, while for German we used gpt-3.5-turbo-0125.
Appendix C Size of Oracle Subsets
In Table 6, the number of clean sentences in the oracle subset for each noisy variant is shown. These oracle subsets are used as performance upper bounds, as explained in Section 4.1.1.
% of all sentences Oracle subset size
NoiseBench split
Clean 100.0 4879
Expert 92.6 4483
Crowd++ 79.4 3786
Crowd 55.3 2554
Distant 59.7 2728
Weak 49.6 2294
LLM 38.2 1705
German split
Clean 100.0 10824
Expert 81.6 8827
LLM 37.8 4095
Table 6: Details about the oracle subset used as an upper performance bound. The table shows the percentage of clean sentences and the absolute number, for each noise type.
Appendix D Baseline for Uniform Noise
Table 7 shows the results from Experiment 1 for uniform noise. We can see the the model is quite robust to uniform noise and that it results in higher test performance, when compared to real noise of the same level. The average difference in F1 scores in 17 percentage points, which is why we focus on the more realistic oracle class-dependent noise simulation method in the main results of Experiments 1 and 2.
Real noise Uniform noise $\Delta$
%Noise F1 %Noise F1 F1
Clean 0 94.0 ±0.0 - - -
Expert 5.5 89.8±0.2 5.4 93.8±0.3 4,0
Crowd++ 15.3 86.7±0.3 16.1 92.8 ±0.5 6.1
Crowd 36.6 70.5±0.6 36.7 88.4 ±0.3 17.9
Distant 31.3 70.8±0.1 31.7 90.0 ±0.5 19.3
Weak 40.4 65.9±0.4 42.2 91.7 ±0.2 25.8
LLM 45.6 62.6±0.4 47.3 89.7 ±0.6 27.1
Average 74.4±0.3 91.40.4 17.0
Table 7: F1 scores on the Clean Test split of the baseline FLERT approach, fine-tuned on different noisy variants of the training set. The scores are averages of 3 runs. The column $\Delta$ (difference) refers to the difference in F1 score on the test split when training on a dataset with real noise compared to uniform noise.
Appendix E Memorization of Crowd, Weak and LLM Noise
Figure 5 is an extension of Section 3.4 and shows the memorization plots from Experiment 2 for the Crowd, Weak and LLM dataset variants. We again observe immediate memorization of real noise and delayed memorization of simulated noise.
Crowd Weak LLM
Real
(a) Real Crowd  - 36.6% noise
(b) Real Weak  - 40.4% noise
(c) Real LLM  - 44.6% noise
Simulated
(d) Simul. Crowd  - 41.3% noise
(e) Simul. Weak  - 41.2% noise
(f) Simul. LLM  - 47.2% noise
Figure 5: Comparison of model performance during extended training, for Crowd, Weak and LLM from NoiseBench. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated noisy label set. The graphs show both the F1 score on the noisy training labels and on the clean training labels, for 3 different noise types. The plots are averages of 3 runs.
Appendix F Additional Experiments on Memorization
In addition to the main Experiment 2, we ran two ablation experiments regarding memorization.
F.1 Effect of Pre-training on Memorization
The first ablation compares fine-tuning a pretrained model and a model with randomly initialized weights. Figure 6 shows this comparison during an extended training run for the Crowd++ training variant, where we used DistilBERT (learning rate of 5e-05). We can see that even without pretraining, the model starts overfitting to the noisy labels and we can observe a large gap between the performance on the clean and noisy labels.
(a) Pre-trained model
(b) Random initialization
Figure 6: Memorization of label noise in DistilBert, using the pretrained model and a model with randomly initialized weights. The experiment was run for one noise type - Crowd++.
F.2 Memorization in a Smaller Model
The second experiment investigates memorization when fine-tuning a smaller model, DistilRoBERTa, because the model we use in the main experiments, XLM-RoBERTa-Large, is quite large. Figure 7 shows an extended training run for three noise types in NoiseBench. We observe the same patterns of immediate and delayed memorization as with the larger model.
Expert Crowd++ Distant
Real
(a) Real Expert  - 5.5% noise
(b) Real Crowd++  - 18% noise
(c) Real Distant  - 31.3% noise
Simulated
(d) Simul. Expert  - 6% noise
(e) Simul. Crowd++  - 15% noise
(f) Simul. Distant  - 39% noise
Figure 7: Comparison of model performance during extended training with a smaller model, Distil-RoBERTa. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated noisy label set. The graphs show the F1 score on the noisy training labels and on the clean training labels, for 3 noise types. The plots are averages of 3 runs.
Appendix G Extended Performance Metrics
In this section we provide extended metrics of the predictive performance of the baseline FLERT method. These metrics and analysis correspond to Experiment 1 from Section 3.3.
G.1 Analysis of Test Errors
We can characterize the model predictions in a similar way as we characterized the different types of errors in Table 1 and Figure 1. Table 8 shows how representative different types of prediction errors are, expressed as a percentage of all errors. We can see that with Expert noise, a majority of the mistakes are wrong entity types. Furthermore, for the Crowd and Distant dataset versions, the largest number of errors is due to missing entities, while for the Weak and LLM datasets, the errors are mostly non-entities or wrong type. This is in line with the characteristics of the noisy datasets themselves, described in Table 1.
For German we make similar observations. For the clean variant, most errors are missing entities. However, for the two noisy variants, which include a large number of noisy non-entity annotations as seen in Table 5, the majority of prediction errors are also non-entity mentions.
% Prediction errors
Missing Non-entity Type Partial
NoiseBench split
Clean 13.9 25.4 29.6 31.1
Expert 12.2 14.4 54.1 19.3
Crowd++ 27.0 11.9 42.1 18.9
Crowd 55.2 6.4 25.3 13.1
Distant 64.6 8.2 14.4 12.8
Weak 14.2 24.5 49.5 11.8
LLM 20.5 34.9 38.3 6.4
German split
Clean 39.9 26.4 16.7 16.9
Expert 17.8 61.9 11.7 8.6
LLM 26.7 61.5 4.8 6.9
Table 8: Overview of the percentage of different types of prediction errors
We also examined the confusion matrices of the predictions. We were able to identify some patterns, regarding which types of errors are more prone to memorization. For most noise types in English NoiseBench, the largest number of prediction mistakes (out of the strings previously seen with a noisy label in the training set) were missing ORG and MISC entities, as well as ORG misclassified as LOC. These mistakes were present in a large number consistently across noise types.
However, we also observed a large number of missing ORG and MISC entities in the predictions when using the clean training set, which indicates that this is an inherently difficult pattern, even when noise is not present. On another hand, the pattern of misclassifying ORG as LOC does not happen when clean data is available. Therefore we can conclude that when this type of noisy pattern is present in the training set, the models are not able to recognize it as noise and are not robust to it.
G.2 Per-Class Metrics
We provide per-class metrics for a more extensive evaluation of the performance of the baseline method, for both German and English, in Table 9.
We can see that both precision and recall for the MISC class are generally lower than the other classes. This is especially noteworthy in the Expert label set of the German split, which does not have a high noise share, but it does have very low performance on MISC. This is however expected, as most of the noisy labels in this label set are related to MISC entities.
LOC ORG PER MISC
Prec. Recall #Ent. Prec. Recall #Ent. Prec. Recall #Ent. Prec. Recall #Ent.
NoiseBench split
Clean 93.9 ±0.6 93.5 ±0.4 1413 93.5 ±0.1 94.6 ±0.1 1909 99.0 ±0.2 99.1 ±0.1 1591 82.9 ±1.0 86.2 ±1.0 812
Expert 81.1 ±0.1 95.5 ±0.6 1413 92.0 ±0.6 81.9 ±0.3 1909 98.4 ±0.7 99.0 ±0.3 1591 85.3 ±0.9 81.4 ±1.1 812
Crowd++ 75.9 ±0.6 94.4 ±0.4 1413 91.8 ±0.3 75.4 ±0.8 1909 97.6 ±0.5 97.8 ±0.3 1591 86.5 ±0.5 70.8 ±0.1 812
Crowd 65.5 ±0.4 90.7 ±1.2 1413 83.7 ±0.8 44.5 ±0.3 1909 93.6 ±1.9 71.7 ±1.2 1591 84.4 ±0.4 47.5 ±0.6 812
Distant 83.8 ±0.7 74.8 ±0.2 1413 85.3 ±1.3 55.9 ±1.0 1909 75.7 ±1.7 84.8 ±0.8 1591 98.6 ±1.4 13.8 ±3.1 812
Weak 52.5 ±0.2 93.1 ±0.1 1413 49.9 ±0.2 34.6 ±0.3 1909 88.5 ±0.3 88.9 ±1.0 1591 84.2 ±0.9 57.3 ±0.9 812
LLM 52.7 ±0.3 84.6 ±0.5 1413 57.7 ±0.9 45.4 ±0.5 1909 95.4 ±0.6 98.3 ±0.1 1591 12.7 ±0.4 11.8 ±0.4 812
German split
Clean 93.0 ±0.7 90.8 ±0.2 1051 80.3 ±0.6 82.3 ±0.3 584 96.6 ±0.4 97.1 ±0.2 1210 77.1 ±1.7 56.5 ±0.6 206
Expert 88.4 ±0.6 83.3 ±0.9 1051 67.1 ±0.8 84.0 ±0.2 584 96.9 ±1.0 96.0 ±0.4 1210 13.1 ±0.3 40.6 ±0.8 206
LLM 63.9 ±0.4 70.3 ±0.9 1051 37.3 ±0.7 82.8 ±1.4 584 66.8 ±0.2 63.7 ±0.6 1210 10.4 ±0.6 18.4 ±1.6 206
Table 9: Per-class metrics of the predictions on the Clean test set